OpenStudy (ttop0816):

PLEASE HELP!!! If 90° < θ < 180° and cotθ = -1/3, then cscθ = _____. √7/3 √10/3 2√2/3 5/3

hartnn (hartnn):

use the pythagorean formula \(\large \csc^2 \theta = 1+ \cot ^2 \theta \)

hartnn (hartnn):

then take square root to get csc theta

OpenStudy (ttop0816):

would it be 2√2/3??

hartnn (hartnn):

1+1/9 10/9 how did u get that ?

OpenStudy (ttop0816):

i think i had a miscalculation!

hartnn (hartnn):

ok, do it again and verify your new answer :)

OpenStudy (ttop0816):

i got squareroot10/3!

hartnn (hartnn):

and you're \(\large \checkmark \) :)

OpenStudy (ttop0816):

@hartnn do you mind checking if my answer is correct for this question?? Position vectors u and v have terminal points of (-6, 3) and (4, -2), respectively. Find the resulting terminal point of 2v - u. (2, -7) (2, 1) (14, -7) (14, 1) <<< i got this!(:

hartnn (hartnn):

14 is correct

hartnn (hartnn):

2(-2) -3 = .. ?

OpenStudy (ttop0816):

omg im so dumb haha thank you again (: im doing a review for trig right now! Could i ask you more questions??

hartnn (hartnn):

sure :)

OpenStudy (ttop0816):

If θ is an angle in standard position that passes through (3, 4), find sin2θ. 2/5 4/5 8/5 24/25 << i got this (:

OpenStudy (ttop0816):

@hartnn

hartnn (hartnn):

2 (3/5) (4/5) = 24/25 you're correct :)

OpenStudy (ttop0816):

horray! You are helping me tons right now! (:

hartnn (hartnn):

always happy to help ^_^

OpenStudy (ttop0816):

how would i solve this?? If cos2 x + 2cosx - 2 = 0, then cosx = _____. -1 ±√3 1 ±√3 ±√3 1

OpenStudy (ttop0816):

@hartnn ! A triangular lot bounded by three streets has a length of 300 feet on one street, 200 feet on the second, and 120 feet on the third. The smallest angle formed by the streets is 30°. Find the area of the lot. 6,000 ft2 9,000 ft2 <<< i got this! 15,000 ft2 30,000 ft2

hartnn (hartnn):

how did u calculate area ?

OpenStudy (ttop0816):

Area = 1/2 * 300 * 75 = 11,250 sq. ft i kinda rounded it to the closest answer;; :P

hartnn (hartnn):

sorry, not so clear with the question... area of triangle with sides, 200, 300, 120 will be 8049.22

OpenStudy (ttop0816):

i know right?? so should i just round it to 9,000ft^2??

hartnn (hartnn):

is there any other info. or diagram with it ? else i would also go with 9000 ...

OpenStudy (ttop0816):

oh okay! how would i solve this?? If cos2 x + 2cosx - 2 = 0, then cosx = _____. -1 plusminus squareroot3 1 plusminus squareroot3 plusminus squareroot3 1

hartnn (hartnn):

is it cos 2x or cos^2 x ?

OpenStudy (ttop0816):

cos^2x!

hartnn (hartnn):

then its a quadratic in cos x answer would be same as x^2+2x-2=0

hartnn (hartnn):

can you solve that quadratic ?

OpenStudy (ttop0816):

x=-1-√3??

OpenStudy (ttop0816):

is the answer -1 plusminus squareroot3 ??

hartnn (hartnn):

yes :)

OpenStudy (ttop0816):

θ is in standard position, and the terminal side passes through the point (3, -5). Find sinθ. -5sqrt31/31 -5sqrt34/34 << ?? -3/5 -4/5